# ANID: How Far Are We? Evaluating the Discrepancies Between AI-synthesized Images and Natural Images through Multimodal Guidance

### v1

@@ C Evaluation Metrics
Aspects Sub_aspects Netrucs Reference
@@
Table 9: The detail of evaluation metrics.
D Human Perceptual Evaluation
Interface for Human Evaluation
The human evaluation of our DNAI dataset was conducted using a custom-designed interface (Figure 11). This interface allowed participants to evaluate AI-generated images in comparison to reference content across three key aspects: Quality, Alignment, and Aesthetics. The reference content included prompts used to generate the images, categorized into three types: 1) text, 2) image, and 3) text and image. The AI-generated images evaluated were produced by four different models: SD_V2.1, SD_XL, Versatile Diffusion, and DALL·E 2.
Evaluation Process
@@ D Human Perceptual Evaluation
E More Evaluations
@@
From Figure 13, we observe notable variability in the differential rates across categories such as ”indoor” and ”sports” among different generative models. Images in the ”indoor” category exhibit significantly higher discrepancies in alignment (CLIP Score) and aesthetics (LAION-AES) compared to ”sports”, indicating that AI-generated sports images are more aligned with natural images. In contrast, generating natural-looking indoor images remains a challenge due to their complex structures and diverse visual features.
Figure 14 presents trends in SSIM and FID metrics across categories and models, where SSIM evaluates pixel-level similarity, and FID measures distribution-level content alignment. The ”accessory” category consistently shows higher FID values, reflecting poorer content distribution, while the ”animal” category achieves higher SSIM scores, demonstrating better pixel-level similarity. These findings suggest that ”accessory” and ”indoor” images pose greater challenges for generative models in maintaining semantic consistency and achieving balanced content generation compared to other categories.
@@
Figure 14: The FID and SSIM of AI-generated images from different source models and categories. The FID and SSIM are used to evaluate image quality at the pixel level and content distribution level.
F Discussion
Safety Mechanism Activation in SD 1.4, 1.5 DALL·E 2
T2I I2I TI2I T2I(%) I2I(%) TI2I(%)
SD v1.4 104 263 141 4.16 10.51 5.64
SD v1.5 133 270 145 5.32 10.79 5.80
DALL·E 2 20 - - 4 - -
Table 10: Unsafe Prompt.
In the rapidly evolving field of AI-generated content (AIGC), the balance between model robustness and safety mechanisms remains a critical area of study. A noteworthy issue has emerged with popular image synthesis models, such as Stable Diffusion v1.4, v1.5, and DALL·E 2, which occasionally fail to generate images in response to seemingly benign prompts. Some examples of these prompts are shown in Figure 10. This unexpected behavior suggests an oversensitivity in the models’ safety filters, which are intended to prevent the creation of inappropriate or sensitive content.

### v2

@@ C.4 Aesthetic
Table 7: The detail of evaluation metrics.
Aspects Sub aspects Metrics Reference
@@
QAlign (aesthetic) FALSE
Figure 12: We evaluate nine image-generating models from 14 aspects. The numerical values in the radar chart represent the mean difference rate of each model and each metric.
Appendix D Evaluation Results
D.1 Naive Quality
D.2 Applicability
We put the applicability results of Image Recognition and Semantic Segmentation in Tables 8 and  9.
Table 8: Image Recognition Discrepancy Rate ($DR_{re}$).
GALIP DFGAN SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
T2I 68.60 85.46 66.14 65.50 68.12 64.88 64.88 65.36 69.26
I2I - - 94.29 94.07 40.90 83.30 27.96 47.98 -
TI2I - - 61.62 59.71 60.05 55.54 27.46 - -
Table 9: IoU of Object Segmentation.
GALIP DFGAN SD14 SD15 VD SD21 SDxl D·E 2 D·E 3
T2I 0.23 0.24 0.26 0.25 0.26 0.25 0.26 0.26 0.28
I2I - - 0.38 0.38 0.69 0.50 0.82 0.42 -
TI2I - - 0.43 0.43 0.29 0.53 0.82 - -
D.3 Human Perceptual Evaluation
D.3.1 Interface for Human Evaluation
The human evaluation of our D-ANI dataset was conducted using a custom-designed interface (Figure 14). This interface allowed participants to evaluate AI-generated images in comparison to reference content across three key aspects: Quality, Alignment, and Aesthetics. The reference content included prompts used to generate the images, categorized into three types: 1) text, 2) image, and 3) text and image. The AI-generated images evaluated were produced by four different models: SD21, SDxl, Versatile Diffusion, and DALL·E 2.
D.3.2 Evaluation Process
@@ D.3.4 Evaluation Metrics
Figure 17: The FID, CLIP-FID and Inception Score (IS) of AI-generated images from different source models and categories from content distribution level.
D.3.5 Summary
D.4 Discrepancies Across Categories
@@
From Figure 16, we observe notable variability in the differential rates across categories such as ‘indoor’ and ‘sports’ among different generative models. Images in the ‘indoor’ category exhibit significantly higher discrepancies in alignment (CLIP Score) and aesthetics (LAION-AES) compared to ‘sports’, indicating that AI-generated sports images are more aligned with natural images. In contrast, generating natural-looking indoor images remains a challenge due to their complex structures and diverse visual features.
Figure 17 presents trends in SSIM, FID, CLIP-FID, and Inception Score. The ‘accessory’ category consistently shows higher FID values, reflecting poorer content distribution, while the ‘animal’ category achieves higher SSIM scores, demonstrating better pixel-level similarity. These findings suggest that ‘accessory’ and ‘indoor’ images pose greater challenges for generative models in maintaining semantic consistency and achieving balanced content generation compared to other categories.
Additionally, the results in Figures 17 and Figures 16 show that compared to other generative models, images generated by D·E 3 display greater differences from natural reference images across various metrics. This might be due to D·E 3’s tendency to produce high-quality images with a surreal style, whereas other models primarily focus on generating images with a more realistic style (see Appendix Sec. B, Figure 9 for visual comparisons).
@@
Appendix E Discussion
Table 10: Prompts trigger the safe-checker of different models.
T2I I2I TI2I T2I(‰) I2I(‰) TI2I(‰)
SD14 104 263 141 4.16 10.51 5.64
SD15 133 270 145 5.32 10.79 5.80
D·E 2 48 - - 9.60 - -
D·E 3 45 - - 9.00 - -
Unexpected Safety Mechanism Activation.
In the rapidly evolving field of AI-generated content (AIGC), the balance between model robustness and safety mechanisms remains a critical area of study. A noteworthy issue has emerged with popular image synthesis models, such as Stable Diffusion v1.4, v1.5, DALL·E 2 and DALL·E 3, which occasionally fail to generate images in response to seemingly benign prompts. Some examples of these prompts are shown in Figure 13. This unexpected behavior suggests an oversensitivity in the models’ safety filters, which are intended to prevent the creation of inappropriate or sensitive content.